# FIRE: A Failure-Adaptive Reinforcement Learning Framework for Edge Computing Migrations

### v2

@@ 3.2 Reward Function
We next specify the reward function of our MDP. We aim to learn a policy that minimizes the total cost of the migration policy, including both operational costs and failure and backup costs.
@@ 3.2.1 Operational Costs
As multiple users share the resources at an edge node, users will experience a compute delay cost. Modelling the process at each AP as an M/M/1 queue [43], the total computing delay experienced across all APs at time $t$ is [12]:
@@
Service profiles are not migrated greedily along with the user, because doing so incurs a service migration cost for the network operator, requiring the operator to balance the communication delay and migration cost. The migration cost $m_{i,j}$, which includes the operational and energy costs on network devices like routers and switches [8], is a function of the distance across two access points $i$ and $j$. Hence, the total migration cost of $K$ users at time $t$ is
@@
where $\mathbbm{1}_{\{l_{s}^{k}(t-1)=i\}}$ is the indicator function which will take the value of $1$ if the service profile was placed at location (access point) $i$ at time $t-1$.
@@ 3.2.2 Failure and Backup Costs
We next specify the costs incurred by provisioning backups in case of failures.
Backup costs. There is a cost $B(t)$ incurred by storing the backup at a server. This cost is a sum of the backup migration cost and the backup storage cost, as it takes up space and prevents other content from being stored.
@@
The failure cost $F$ is experienced at time $t$ when 1) location $l_{s}^{k}(t)$ has a failure at time $t$ $(f^{k}_{ind}(t)=1)$, and b) there is no backup in the system ($b_{u}^{k}(t)=\text{no backup}$), or when the placement of the backup $b_{u}^{k}(t)$ equals the location of the main service $l_{s}^{k}(t)$. Otherwise, if there is a backup placed, there will be no failure cost experience, and the cost will be much lower. Note, however, that the presence of failures themselves, i.e., $f_{ind}(t)$, is independent of the user actions.
4 Service Migration of Shared Service Profiles in Light of Rare Events
Figure 2: Multiple users share a service profile for their task application, e.g. a common game environment or neural network. For instance, the users in red share the service profile in red, which has multiple copies in the network.
@@ 4 Service Migration of Shared Service Profiles in Light of Rare Events
where $d^{comm}_{i,\hat{i}}$ is the communication delay from location $i$ to $\hat{i}$, same as Section 3. The users at location $i$ access the nearest AP which contains the SP type they require ($\min_{\hat{i}\in\{\hat{i}|l^{SPj}_{\text{loc}\ \hat{i}}=1\}}d^{\text{comm}}_{i,\hat{i}}$). User QoS is also affected by the total computing delay experienced across all APs at time $t$, which is [12]:
@@
The total migration cost at time $t$ of required migration is
@@ 5 Cost Minimization in Light of Rare Events
where $w_{D},w_{C},$ etc. are the corresponding weights, which can be adjusted based on the priority of different costs. The decision variable $\pi(s,a)=\{P_{r}(a_{t}=a|s_{t}=s)\}$ is the network operator’s policy, the probability it will take each action $a(t)$ (service placement and backup storage decision for all users) given the state $s(t)$. The first constraint indicates that the state dynamics follows the transition probability of the system in Eq. (7), incorporating rare event transitions, while the other constraints indicate the delay, migration, backup and failure costs respectively. Next, we introduce the optimization problem for the scenario where users share service profiles; note that we aim to minimize the same types of costs as in Eq. (18) with individual service profiles.
@@
Deciding a migration path for the user across $T$ timeslots is similar to the shortest path problem. However, the network operator may not have information on mobility distributions and the corresponding latency costs, complicating look-ahead solutions. Since the costs of different users are coupled through service profiles’ sharing of AP resources and consequent computing delays (see Eqs. (9) and (16)) , finding the optimal migration strategies becomes yet more difficult. Thus, Markov Decision Processes and Reinforcement Learning have been used to derive service migration policies in the edge computing literature [11, 16, 17]. The difficulty in our problem further arises as we model occasional server failures, which impact the smooth and safe functioning of applications. As these failures occur at a low probability, and because reinforcement learning relies on past reward data, the learner may miss the importance of these rare events and be unable to learn an optimal policy adaptive to server failures. We introduce our solution in the next section.
@@ 6 Reinforcement learning in the presence of Server Failures
In this section, we present FIRE: Failure-adaptive Importance sampling for Rare Events, a reinforcement learning based framework in an edge computing digital twin, to optimize service migration and backup placement in light of potential server failures. We define rare events formally in Subsection 6.1, introduce importance sampling in Subsection LABEL:subsec:impSamp, present our algorithm (ImRE) and its proof in Subsections 6.2 and 6.3 respectively, and present an eligibility traces version of the algorithm (ETAA) for faster convergence in the Appendix. We propose a Deep Q learning and Actor Critic version of our algorithm in Section 7 to accommodate large state and action spaces, e.g., for many servers or users across a long amount of time.
6.1 Rare Events And Their Impact On The Value Function
@@
The Q-value (action-value) of a state action pair $(s,a)$ is the expected-return of taking action $a$ at state $s$, following $\pi$.
$Q^{\pi}(s,a)=\mathbb{E}_{\pi}[\sum_{k=1}^{\infty}\gamma^{k}r(s(t),a(t),s(t+1))|s_{t}=s,a_{t}=a].$ (24)
$Q^{\pi}(s,a)=\sum_{s^{\prime}\in S}p(s^{\prime}|s,a)[r(s,a,s^{\prime})+\gamma\sum_{a^{\prime}\in A}\pi(a^{\prime}|s^{\prime})Q^{\pi}(s^{\prime},a^{\prime})]$ (25)
@@
6.2 ImRE: Q-Learning in Light of Rare Events
@@
When $q$ is well crafted, variance in the estimates will be reduced. The use of importance sampling in temporal difference (TD) based RL was proposed in [39, 20], where the state transition probabilities followed distribution $q(s^{\prime}|s,a)$ rather than the true distribution $p(s^{\prime}|s,a)$. For our service migration and backup placement problem, we will sample the rare events at probability $\hat{\epsilon}(s)$ (analogous to $q(s^{\prime}|s,a)$), a higher rate than $\epsilon(s)$ (analogous to $p(s^{\prime}|s,a)$), which is the probability of visiting a rare event state from state $s$. This allows us to sample rare events at a rate proportional to their contribution to the value of state $s$. Unlike [39, 20]’s importance sampling based value estimation given a fixed policy, our aim is to iterate over policies and learn an optimal policy $\pi^{*}$ through importance sampling based Q-learning.
Therefore, to obtain $\hat{\epsilon}(s)$, we define $T(s,a)$ as the contribution of the rare states $s^{\prime}\in T$ towards the Q-value of the state action pair $(s,a)$, and $U(s,a)$ as the contribution of the normal states $s^{\prime}\in S\backslash T$ towards the Q-value of the state action pair $(s,a)$:
@@
Nevertheless, updating at the state action $(s,a)$ level may lead to slower convergence, especially with an increasing number of access points, leading to an exponential increase in the state-action combinations. For example, with 9 access points, there will be 360 states and 90 actions, leading to 32400 state-action combinations. We can reduce this complexity by exploiting our earlier observation that the probability of rare events (i.e., server failures) happening is independent of the actions chosen. Therefore, we propose updating $T$ and $U$ (the contributions of the rare events and non-rare events respectively towards the Q-value) at the state level:
@@
Algorithm description: The algorithm is presented in Algorithm 1 (ImRE). Firstly, we initialise $\hat{Q}(s,a)$, $\hat{T}(s)$ and $\hat{U}(s)$ to be 00 (for all states), and $\hat{\epsilon}(s)$ to be $\frac{1}{2}$ for all states. We also initialise the learning rates $\alpha^{t},\alpha^{t}_{T},\alpha^{t}_{U}$. For every timeslot, the importance sampling rare event probability $\hat{\epsilon}(s)$ will determine whether or not a rare-event (failure) occurs. Thereafter, the rest of the state transition will occur according to the probability distribution $h(\tilde{s}|s,a)$, where $\tilde{s}(t)=(l_{u}(t),l_{s}(t),b_{ind}(t))$. This would result in a new state $s^{t+1}$, and a reward value $r^{t+1}=r(s^{t},a^{t},s^{t+1})$ (line 4). Based on the next state $s^{t+1}$, a new action is selected according to the $\beta$-greedy policy (lines 5-6), which is the same as the traditional $\epsilon$-greedy policy in Q-learning, where with probability $\beta$ the greedy action is selected and with probability $1-\beta$ a random action is selected, for exploration. We call it $\beta$-greedy to avoid confusion with our rare event probability $\epsilon$.
@@
Algorithm 1 ImRE: Importance Sampling Q-Learning for Rare Events
@@
6.3 Proofs of Boundedness and Convergence to Optimality
In this subsection, we prove the convergence of ImRE, our importance sampling based Q-learning algorithm (Algorithm 1).
Firstly, we show that the sequence of updates in Algorithm 1 is bounded, in the following theorem. We provide a summarized version of the proof here, with its full version in our Technical Report [45]. The main idea of the proof is as follows: We rewrite Algorithm 1’s update equation, through adding and subtracting terms. We rewrite the equation in the form of $\hat{Q}^{t+1}(s,a)=\hat{Q}^{t}(s,a)+\alpha^{t}[F_{s,a}(\hat{Q}^{t})-\hat{Q}^{t}(s^{t},a^{t})+M_{s}(t)]$. Invoking Theorem 1 of [46] on stochastic approximation algorithms, we finally prove convergence by showing that the conditions for the sequence $F_{s,a}(\hat{Q})$ to be bounded are met.
@@ 6.3 Proofs of Boundedness and Convergence to Optimality
For stepsizes following the conditions $\sum_{t}\alpha(t)=\infty$ and $\sum_{t}\alpha^{2}(t)\leq\infty$, and discount factor $\gamma\in(0,1)$, the sequence of the $\hat{Q}$ updates in Algorithm 1 is bounded, with probability 1.
@@
Algorithm 1’s update equation is
$\hat{Q}(s^{t},a^{t})\leftarrow\hat{Q}(s^{t},a^{t})+\alpha[w_{t}(r^{t+1}+\gamma\hat{Q}(s^{t+1},a^{*}))-\hat{Q}(s^{t},a^{t})],$ (35)
where $a^{*}=\text{argmax}_{b}\hat{Q}(s^{t+1},b)$.
We rewrite Algorithm 1’s update equation through adding and subtracting the term $\sum_{s^{\prime}\in S}\hat{p}(s^{\prime}|s^{t},a^{t})(r^{t}(s^{t},a^{t},s^{\prime})+\gamma\max_{y}\hat{Q}^{t}(s^{\prime},y))$.
$\hat{Q}^{t+1}(s,a)$ can be further rewritten as
$\hat{Q}^{t+1}(s,a)=\hat{Q}^{t}(s,a)+\alpha^{t}[F_{s,a}(\hat{Q}^{t})-\hat{Q}^{t}(s^{t},a^{t})+M_{s}(t)],$
for all $(s,a)$, where $F_{s,a}=[F_{s1,a1},...,F_{|S\times A|}]^{T}$ is defined by
$\begin{split}&F_{s,a}(\hat{Q}^{t}(s,a))\\ &=\sum_{s^{\prime}\in S}\hat{p}(s^{\prime}|s,a)(r^{t}(s,a,s^{\prime})+\gamma\max_{y}\hat{Q}^{t}(s^{\prime},y)),\end{split}$
and $\{M(t)\},M(t)=[M_{1,1}(t),M_{1,2}(t),...,M_{|S\times A|}(t)]$ is a sequence of random vectors in $\mathbb{R}^{|S\times A|}$ satisfying
$\mathbb{E}[M_{m}(t)|m<t]=0$
$||M(t)||\leq K(1+||\hat{Q}^{t}||)$
for some constant $K>0$. $||M(t)||$ is bounded because the probabilities $\epsilon(s^{t})$ and $1-\epsilon(s^{t})$ are fixed, while $\hat{\epsilon}(s^{t})$ and hence $1-\hat{\epsilon}(s^{t})$ are bounded in our algorithm (line 15). This results in the importance weight $w^{t}$ (Eq. (33)) being bounded.
We show that the property $||F(\hat{Q})||_{\infty}\leq b||\hat{Q}||_{\infty}+d$ holds, where $||\quad||_{\infty}$ is the supremum norm.
$||F(\hat{Q})||_{\infty}$
$=\max_{i,a}|\sum_{s^{\prime}\in S}\hat{p}(s^{\prime}|i,a)(r(i,a,s^{\prime})+\gamma\max_{y}\hat{Q}(s^{\prime},y))|$ (36)
$\leq\max_{i,a}[|\sum_{s^{\prime}\in S}\hat{p}(s^{\prime}|i,a)r(i,a,s^{\prime})|$
$+|\sum_{s^{\prime}\in S}\hat{p}(s^{\prime}|i,a)\gamma\max_{y}\hat{Q}(s^{\prime},y)|]$ (37)
$\leq\max_{i,a}|\sum_{s^{\prime}\in S}\hat{p}(s^{\prime}|i,a)r(i,a,s^{\prime})|$
$+\gamma\max_{i,a}|\sum_{s^{\prime}\in S}\hat{p}(s^{\prime}|i,a)\max_{z,y}\hat{Q}(z,y)|$ (38)
$=\max_{i,a}|\sum_{s^{\prime}\in S}\hat{p}(s^{\prime}|i,a)r(i,a,s^{\prime})|+\gamma\max_{i,a}\sum_{s^{\prime}\in S}\hat{p}(s^{\prime}|i,a)||\hat{Q}||_{\infty}$
$=b||\hat{Q}||_{\infty}+d$
where $b=\gamma\max_{i,a}\sum_{s^{\prime}\in S}\hat{p}(s^{\prime}|i,a)=\gamma<1$ and $d=\max_{i,a}|\sum_{s^{\prime}\in S}\hat{p}(s^{\prime}|i,a)r(i,a,s^{\prime})|=\max_{i,a}|r(i,a,s^{\prime})|$. The first inequality (from Eq. (36) to (37)) is true by the triangle inequality. Therefore, by Theorem 1 of [46] on stochastic approximation algorithms, the sequence of $\hat{Q}$ updates in Algorithm 1 is bounded, with probability 1. ∎
Next, to help us prove that our algorithm ImRE converges to optimality, we invoke the following theorem from [47]:
@@
The random process $\{\triangle_{t}\}$ taking values in $\mathbb{R}^{n}$ and defined as
$\triangle_{t+1}(x)=(1-\alpha_{t}(x))\triangle_{t}(x)+\alpha_{t}F_{t}(x)$ (39)
converges to zero with probability 1 when the following assumptions hold:
[…]
In the following theorem, we show that our algorithm converges, and that it converges to the optimal policy. We provide a summarized version of the proof here, with its full version in our Technical Report [45]. The main idea of the proof is as follows. Firstly, we show that the optimal policy of the true system is a fixed point of the following equation with our simulator’s transition probability $\hat{p}(s^{\prime}|s,a)$ and the importance sampling correction weight $w_{t}(s)$: $Q^{*}(s,a)=\sum_{s^{\prime}\in S}\hat{p}(s^{\prime}|s,a)w_{t}(s)[r(s,a,s^{\prime})+\gamma\max_{b}Q^{*}(s^{\prime},b)].$ We define the operator $\mathbf{H}\hat{Q}=\sum_{s^{\prime}\in S}\hat{p}(s^{\prime}|s,a)w_{t}(s)[r(s,a,s^{\prime})+\gamma\max_{b}\hat{Q}(s^{\prime},b)]$, and show that $\mathbf{H}\hat{Q}$ is a contraction mapping. Next, we invoke the result of Corollary 3 on the convergence of random processes. We rewrite the update equation of Algorithm 1, such that it fits the form of Eq. (39) in Corollary 3, and show that the conditions of Corollary 3 hold, using the fact that $\mathbf{H}\hat{Q}$ is a contraction mapping.
@@
If the MDP underlying the reinforcement learning environment is unichain for $\epsilon\in(\delta,1-\delta)$,111$(\delta,1-\delta)$ is the range of values which $\hat{\epsilon}$ takes, as defined in Algorithm 1. If the MDP defined by transition probability $p(s^{\prime}|s,a)$ is unichain for one value in $(\delta,1-\delta)$, it is unichain for all values [20]. for stepsizes following the conditions $0\leq\alpha_{t}\leq 1$, $\sum_{t}\alpha_{i}(t)=\infty$ and $\sum_{t}\alpha_{i}^{2}(t)\leq\infty$, and discount factor $\gamma\in(0,1)$, Algorithm 1 converges to the optimal policy $Q^{*}$.

### v3

@@ 3.2 Reward Function
We next specify the reward function of our MDP. We aim to learn a policy that minimizes the total cost of migration, including operational, failure and backup costs.
@@ 3.2.1 Operational Costs
As multiple users share the resources at an edge node, users will experience a compute delay. Modelling the process at each AP as an M/M/1 queue, the total computing delay experienced across all APs at time $t$ is [8]:
@@
The migration cost $m_{i,j}$, which includes the operational and energy costs on network devices like routers and switches [5], is a function of the distance across two APs $i$ and $j$. The total migration cost of $K$ users at time $t$ is
@@
where $\mathbbm{1}_{\{l_{s}^{k}(t-1)=i\}}$ is the indicator function that equals $1$ if the service profile was placed at location (AP) $i$ at time $t-1$.
@@ 3.2.2 Failure and Backup Costs
We next specify the costs incurred by provisioning backups in case of failures. There is a backup cost $B(t)$ incurred by storing the backup at a server: a sum of the backup migration cost and the backup storage cost, as it takes up space and prevents other content from being stored.
@@
This cost is experienced at time $t$ when i) location $l_{s}^{k}(t)$ has a failure at time $t$ $(f^{k}_{ind}(t)=1)$, and ii) there is no backup in the system ($b_{u}^{k}(t)=\text{no backup}$), or when the placement of the backup $b_{u}^{k}(t)=l_{s}^{k}(t)$, the location of the main service. Otherwise, if there is a backup placed, there will be no failure cost. The presence of failures themselves, i.e., $f_{ind}(t)$, is however independent of the user actions.
4 Migration of Shared Service Profiles
Figure 2: Multiple users share a service profile (SP) for their task, e.g. a common game environment or neural network. For instance, the users in red share the SP in red, which has multiple copies in the network.
@@ 4 Migration of Shared Service Profiles
where $d^{comm}_{i,\hat{i}}$ is the communication delay from location $i$ to $\hat{i}$, same as Section 3. The users at location $i$ access the nearest AP which contains the SP type they require ($\min_{\hat{i}\in\{\hat{i}|l^{SPj}_{\text{loc}\ \hat{i}}=1\}}d^{\text{comm}}_{i,\hat{i}}$). QoS is also affected by the total computing delay experienced across all APs at time $t$ [8]:
@@
The total migration cost at time $t$ is
@@ 5 Cost Minimization in Light of Rare Events
where $w_{D},w_{C},$ etc. are the corresponding weights, which can be adjusted based on the priority of different costs. The decision variable $\pi(s,a)=\{P_{r}(a_{t}=a|s_{t}=s)\}$ is the network operator’s policy, the probability it will take each action $a(t)$ (service placement and backup storage decision for all users) given the state $s(t)$. The first constraint indicates that the state dynamics follows the transition probability of the system in Eq. (7), incorporating rare event transitions, while the other constraints indicate the delay, migration, backup and failure costs respectively. Next, we introduce the optimization problem for the shared service profiles scenario; note that we aim to minimize the same types of costs as in Eq. (18) with individual service profiles.
@@
Deciding a migration path for the user across $T$ timeslots is similar to the shortest path problem. However, the network operator may not have information on mobility distributions and the corresponding latency costs, complicating look-ahead solutions. Furthermore, the costs (e.g. computing delays, Eqs. (9) and (16)) of different users are coupled through service profiles’ sharing of AP resources. Thus, Markov Decision Processes and Reinforcement Learning have been used to derive service migration policies [7, 10, 13]. The difficulty in our problem further arises as we model occasional server failures. As these failures occur at a low probability, and because reinforcement learning relies on past reward data, the learner might overlook rare events and fail to learn an optimal failure-adaptive policy. We introduce our solution in the next section.
@@ 6 Reinforcement learning in the presence of Server Failures
In this section, we present FIRE: Failure-adaptive Importance sampling for Rare Events, a reinforcement learning based framework in an edge computing digital twin. While [16] and [14] also use importance sampling to sample rare events, these two works involve estimating the value function given a fixed policy. Their proof techniques also do not generalize to the $Q$-learning. In contrast, our algorithms involve learning the optimal policy $\pi^{*}$, which involves policy changes. Based on our algorithm, we propose novel boundedness and convergence proofs. We present our algorithm (FIRE-ImRE) and its proof in Subsections 6.2 and 6.3 respectively, and present an eligibility traces version (ETAA) for faster convergence in the Appendix. We propose a Deep Q learning and Actor Critic version of our algorithm in Section 7 to accommodate large state and action spaces. FIRE is compatible with any reinforcement learning formulation, and our algorithms are examples of this.
6.1 Rare Events and the Value Function
@@
The Q-value of a state action pair $(s,a)$ is the expected-return of taking action $a$ at state $s$, following $\pi$.
$Q^{\pi}(s,a)=\mathbb{E}_{\pi}[\sum_{k=1}^{\infty}\gamma^{k}r(s(t),a(t),s(t+1))|s_{t}=s,a_{t}=a]$
$=\sum_{s^{\prime}\in S}p(s^{\prime}|s,a)[r(s,a,s^{\prime})+\gamma\sum_{a^{\prime}\in A}\pi(a^{\prime}|s^{\prime})Q^{\pi}(s^{\prime},a^{\prime})]$
@@
6.2 FIRE-ImRE: Q-Learning in Light of Rare Events
@@
When $q$ is well crafted, variance in the estimates will be reduced. For our service migration and backup placement problem, we will sample the rare events at probability $\hat{\epsilon}(s)$ (analogous to $q(x)$), a higher rate than $\epsilon(s)$ (analogous to $p(x)$), which is the probability of visiting a rare event state from state $s$. Therefore, to obtain $\hat{\epsilon}(s)$, we define $T(s,a)$ as the contribution of the rare states $s^{\prime}\in T$ towards the Q-value of the state action pair $(s,a)$, and $U(s,a)$ as the contribution of the normal states $s^{\prime}\in S\backslash T$ towards the Q-value of the state action pair $(s,a)$:
@@
Updating at the state action $(s,a)$ level may lead to slower convergence, especially with an increasing number of access points. For example, with 9 access points, there will be 360 states and 90 actions, leading to 32400 state-action combinations. We can reduce this complexity by exploiting our earlier observation that the probability of rare events (i.e., server failures) happening is independent of the actions chosen. Therefore, we propose updating $T$ and $U$ (the contributions of the rare events and non-rare events respectively towards the Q-value) at the state level:
@@
Algorithm description: The algorithm is presented in Algorithm 1 (FIRE-ImRE). Firstly, we initialise $\hat{Q}(s,a)$, $\hat{T}(s)$ and $\hat{U}(s)=0$, and $\hat{\epsilon}(s)=\frac{1}{2}$, for all states, and initialise the learning rates $\alpha^{t},\alpha^{t}_{T},\alpha^{t}_{U}$. For every timeslot, the importance sampling rare event probability $\hat{\epsilon}(s)$ will determine whether or not a rare-event (failure) occurs. Thereafter, the rest of the state transition will occur according to the probability distribution $h(\tilde{s}|s,a)$, where $\tilde{s}(t)=(l_{u}(t),l_{s}(t),b_{ind}(t))$. This would result in a new state $s^{t+1}$, and a reward value $r^{t+1}=r(s^{t},a^{t},s^{t+1})$ (line 4). Based on the next state $s^{t+1}$, a new action is selected according to the $\beta$-greedy policy (lines 5-6), which is the same as the traditional $\epsilon$-greedy policy in Q-learning, where with probability $\beta$ the greedy action is selected and with probability $1-\beta$ a random action is selected, for exploration. We call it $\beta$-greedy to avoid confusion with our rare event probability $\epsilon$.
@@
Algorithm 1 FIRE-ImRE: Importance Sampling Q-Learning for Rare Events
@@
6.3 Boundedness and Convergence Properties
In this subsection, we prove the convergence of FIRE-ImRE, our importance sampling based Q-learning algorithm (Algorithm 1).
Firstly, we show that the sequence of updates in Algorithm 1 is bounded, in the following theorem.
@@ 6.3 Boundedness and Convergence Properties
For stepsizes that satisfy $\sum_{t}\alpha(t)=\infty$ and $\sum_{t}\alpha^{2}(t)\leq\infty$, and discount factor $\gamma\in(0,1)$, the sequence of the $\hat{Q}$ updates in Algorithm 1 is bounded, with probability 1.
@@
See our Technical Report [32].
The main idea of the proof is as follows: We rewrite Algorithm 1’s update equation, through adding and subtracting terms. We rewrite the equation in the form of $\hat{Q}^{t+1}(s,a)=\hat{Q}^{t}(s,a)+\alpha^{t}[F_{s,a}(\hat{Q}^{t})-\hat{Q}^{t}(s^{t},a^{t})+M_{s}(t)]$. Invoking Theorem 1 of [33] on stochastic approximation algorithms, we finally prove convergence by showing that the conditions for the sequence $F_{s,a}(\hat{Q})$ to be bounded are met. We do this by showing that $||M(t)||$ is bounded, and that the property $||F(\hat{Q})||_{\infty}\leq b||\hat{Q}||_{\infty}+d$ holds, where $||\quad||_{\infty}$ is the supremum norm.
Next, to help us prove that our algorithm FIRE-ImRE converges to optimality, we invoke the following corollary from [34]:
@@
The random process $\{\triangle_{t}\}$ taking values in $\mathbb{R}^{n}$ and defined as $\triangle_{t+1}(x)=(1-\alpha_{t}(x))\triangle_{t}(x)+\alpha_{t}F_{t}(x)$ converges to zero with probability 1 when the following assumptions hold:
$0\leq\alpha_{t}\leq 1,\sum_{t}\alpha_{t}(x)=\infty\ \text{and}\sum_{t}\alpha^{2}_{t}(x)<\infty.$
$||\mathbb{E}[F_{t}(x)]||\leq\gamma||\triangle_{t}||,\text{where}\ \gamma<1.$
$var[F_{t}(x)]\leq C(1+||\triangle_{t}||^{2}),\ \text{for}\ C>0.$
In the following theorem, we show that our algorithm converges, and that it converges to the optimal policy.
@@
If the MDP underlying the RL environment is unichain for $\epsilon\in(\delta,1-\delta)$,111$(\delta,1-\delta)$ is the range of values which $\hat{\epsilon}$ takes, as defined in Algorithm 1. If the MDP defined by transition probability $p(s^{\prime}|s,a)$ is unichain for one value in $(\delta,1-\delta)$, it is unichain for all values [14]. for stepsizes satisfying $0\leq\alpha_{t}\leq 1$, $\sum_{t}\alpha_{i}(t)=\infty$ and $\sum_{t}\alpha_{i}^{2}(t)\leq\infty$, and discount factor $\gamma\in(0,1)$, Algorithm 1 converges to the optimal policy $Q^{*}$.